y varies directly as x and inversely as the square of z. y=24 when x=12 and z=2. Find y when x=64 and z=4.
y=

thank you so much!

Answer by richwmiller(17219)

y=kx/z^2
24=2k/4
24=k/2
k=48
y=48x/z^2
y=48*64/16
y=192